### v1

@@ 3 Inner-interface spatial discretisation
This section explores two new approaches to accelerate AMG by making it compatible with SpMM. Unlike LRCFSAI($k$) and LRCAMG($k$), the preconditioners developed here rely on a different ordering of unknowns, henceforth denoted as inner-interface ordering. To define it, let us recall the single symmetry case of Fig. 2(a) and classify all the grid points into interface: those coupled with other subdomain unknowns, and inner: those that are not. Then, the inner-interface ordering arises from applying the mirrored ordering of Fig. 1 to the inner unknowns first and the interface ones afterwards. Remarkably enough, AMGR, the preconditioner derived in Section 3.2, relaxes this constraint, allowing any ordering of the interface unknowns. Additionally, boundary conditions are no longer required to be symmetric, therefore being compatible with repeated geometries like the illustrative Fig. 2(b).
@@
where $\bar{K}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{inn}}}$, $\bar{B}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{ifc}}}$, and $\bar{C}\in\mathbb{R}^{n_{\text{ifc}}\times n_{\text{ifc}}}$. Additionally, thanks to the mirrored ordering imposed on the inner and interface grid points, we have that $\bar{K}=\mathbb{I}_{2^{s}}\otimes K$ and $\bar{B}=\mathbb{I}_{2^{s}}\otimes B$. On the other hand, $C$ satisfies Eq. 6, and consequently, SpMM can be exploited to accelerate all the submatrix products while reducing their setup costs and memory footprint.
@@ 3.1 Schur complement-based AMG
Given the poor results obtained with LRCAMG($k$), our aim now is to exploit the structure of $\bar{K}$, $\bar{B}$ and $\bar{C}$ to accelerate the AMG approximation of $A$, $M_{A}\simeq A^{-1}$. By virtue of Eq. 36, the inverse of $A$’s LDU factorisation reads:
@@
where $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}$ is the Schur complement of $\bar{K}$. Then, we can derive valid preconditioners by seeking approximations of $\bar{K}^{-1}$ and $\bar{S}^{-1}$. In our case, given that $\bar{K}^{-1}=\mathbb{I}_{2^{s}}\otimes K^{-1}$, we will approximate $\bar{K}^{-1}$ through the AMG of $K$, $M_{K}\simeq K^{-1}$. This has two major advantages: on the one hand, the setup costs and memory footprint of $M_{K}$ are $2^{s}$ times smaller than those of $M_{\bar{K}}$. On the other, its application is compatible with SpMM.
Regarding the approximation of $\bar{S}^{-1}$, the better it is, the closer the action of our Schur-based AMG will be to that of $M_{A}$. Given the favourable results of Section 2.1, we will build an approximation of $\bar{S}^{-1}$ based on LRCFSAI($k$). For simplicity, let us assume a single reflection symmetry. As before, applying a mirrored ordering to the interface grid points ensures that $\bar{C}$ satisfies the following structure:
@@
where $C_{1},C_{2}\in\mathbb{R}^{n_{\text{ifc}}/2\times n_{\text{ifc}}/2}$ account for the interface-interface couplings of the base mesh with itself and its mirroring, respectively. On the other hand, we have that:
(39) $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}=\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}K^{-1}B=\begin{pmatrix}S_{1}&C_{2}\\ C_{2}&S_{1}\end{pmatrix},$
@@
The fact that $\bar{S}$ contains $K^{-1}$ prevents us from using AMG to approximate $\bar{S}^{-1}$. However, given the excellent results obtained with LRCFSAI($k$), it makes sense to mimic its strategy. In particular, we will first build a rough approximation of $\bar{S}^{-1}$ based on an FSAI of $S_{1}$ and compatible with SpMM. Then, we will correct it by means of low-rank corrections to make it approach the (block diagonal version of the) most accurate Schur complement at hand:
(41) $\bar{S}_{\text{AMG}}\coloneqq\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}M_{K}B,$
@@
With this aim, let $L_{K}$ be the lower Cholesky factor of $K$, and $G_{K}\simeq L_{K}^{-1}$ its lower FSAI factor. Then, we have that $G_{K}^{T}G_{K}\simeq K^{-1}$, and it is meaningful to approximate $S_{1}^{-1}$ through an FSAI of $C_{1}-B^{T}G_{K}^{T}G_{K}B\simeq S_{1}$:
(42) $S_{1}^{-1}\simeq G_{S_{1}}^{T}G_{S_{1}}.$
Finally, by recalling Eq. 40 and the fact that $S_{1}$ is substantially denser than $C_{2}$, we can build the sought approximation of $\hat{S}^{-1}$ as follows:
(43) $\hat{S}^{-1}\simeq\mathbb{I}_{2}\otimes G_{S_{1}}^{T}G_{S_{1}},$
@@
Regardless of how we approximate $\bar{S}^{-1}$, Eq. 37 would require applying $M_{K}$ twice, whereas $M_{A}$ is only applied once. In order to avoid such an expensive overhead, we can define our Schur-based AMG, henceforth denoted AMGS, as follows:
(45) $M_{\text{Schur}}\coloneqq\begin{pmatrix}\mathbb{I}&-G_{\bar{K}}^{T}G_{\bar{K}}\bar{B}\\ &\mathbb{I}\end{pmatrix}\begin{pmatrix}\text{AMG}_{\bar{K}}&\\ &Q_{s}^{-1}\hat{S}_{\text{LRC}}^{-1}Q_{s}\end{pmatrix}\begin{pmatrix}\mathbb{I}&\\ -\bar{B}^{T}M_{\bar{K}}&\mathbb{I}\end{pmatrix},$
where we compacted the notation of $M_{\bar{K}}\coloneqq\mathbb{I}_{2^{s}}\otimes M_{K}$ and replaced its extra application with the FSAI of $\bar{K}$, also compacted to $G_{\bar{K}}^{T}G_{\bar{K}}\coloneqq\mathbb{I}_{2^{s}}\otimes G_{K}^{T}G_{K}$. Remarkably enough, this substitution did not deteriorate the convergence of AMGS significantly, but made it nonsymmetric and, therefore, incompatible with PCG.
The main advantages of AMGS are the fact that its matrix multiplications are compatible with SpMM, together with the smaller memory footprint and setup costs that this allows. Table 3 summarises the results obtained with AMGS on the model problem of Eq. 25. As shown, the combination of FSAI and low-rank corrections was not enough to substantially improve its performance with respect to LRCAMG($k$). Consequently, we discarded its parallel implementation and developed the better alternative of Section 3.2.
@@
AMGS 6 21 23 22
@@ 3.2 Multigrid reduction
As the model problem of Eq. 25 confirmed, the standard AMG algorithm cannot take advantage of spatial symmetries and needs to be appropriately adapted. In particular, we will develop an AMG reduction framework, hereafter denoted as AMGR, that leverages reflection, translational and rotational symmetries regardless of the boundary conditions. Remarkably enough, AMGR only uses SpMM on the inner unknowns, therefore allowing any ordering of the interface. It relies on a prolongation, $P$, with a block structure similar to Eq. 36:
(46) $P=\begin{pmatrix}\bar{W}\\ \mathbb{I}_{n_{c}}\end{pmatrix}\in\mathbb{R}^{n\times n_{c}},$
where $\bar{W}\in\mathbb{R}^{n_{f}\times n_{c}}$ and $\mathbb{I}_{n_{c}}\in\mathbb{R}^{n_{c}\times n_{c}}$. Additionally, $n_{f}$ and $n_{c}$ correspond to the number of fine and coarse nodes as defined by the multigrid reduction.
@@
Ideally, to have the fastest possible coarsening, we should only identify as coarse the interface unknowns, i.e., the $n_{\text{ifc}}$ unknowns associated with the interface block, $\bar{C}$. However, realistically sized problems make it impossible to accurately interpolate inner unknowns solely using interface values. This is due to the large connection distance that may occur. Indeed, if there are $n_{\text{inn}}$ inner unknowns, the maximum inner-interface distance will be of the order of $d]{n_{\text{inn}}}$, where $d$ stands for the geometrical dimension of the problem. Then, a 3D problem with about a million inner unknowns would result in distances of about 100 units, which exceeds the applicability of long-distance interpolation formulas such as Extended+I (ExtI) [18] or dynamic-pattern Least Squares Fit (LSF) [43].
@@
Hence, to allow for an accurate interpolation, we need to convert some inner nodes into coarse. In principle, we can apply any standard coarsening strategy to the inner part of each subdomain’s sub-block. That is, by choosing a strength of connection measure, e.g., a classical strength measure, filtering the resulting adjacency graph to obtain $T$, and running a maximum independent set (MIS) algorithm on $T$ to select which inner nodes become coarse. However, even if many connections are retained in $T$, the number of coarse nodes generally becomes too high, making the resulting operator’s complexity impractical. To cope with this issue, we use an aggressive coarsening strategy designed ad hoc for this problem. Instead of selecting independent nodes from $T$, we select independent nodes from $T^{k}$, for a small power $k$. Using $T^{k}$ is equivalent to constructing an independent set by considering dependent to each other all nodes at a distance smaller or equal to $k$. Larger distances between coarse nodes are not an issue if long-distance interpolations are used. ExtI interpolation is most effective when $k$ is limited to 2. Conversely, LSF allows for larger values of $k$ and coarser operators.
Table 4 summarises the results obtained with AMGR on the model problem of Eq. 25. Adopting the aggressive coarsening that symmetries and repeated geometries induce does not harm convergence and allows for a small and sparse coarse level operator, $A_{c}=P^{T}AP$. Then, if $A_{c}$ is not small enough, we can approximate its inverse using a standard AMG. Additional advantages of AMGR include the compatibility with SpMM of the FSAI used as $\bar{K}$-smoother, reducing its memory footprint and setup costs.
@@
Given the promising results obtained with AMGR, we tackled its parallel implementation, which is discussed in Section 4, and tested it in industrial CFD applications in Section 5.

### v2

@@ 3 Inner-interface spatial discretisation
This section explores two new approaches to accelerate AMG by making it compatible with SpMM. Unlike LRCFSAI($k$) and LRCAMG($k$), the preconditioners developed here rely on a different ordering of unknowns, henceforth denoted as inner-interface ordering. To define it, let us recall the single symmetry case of Fig. 2(a) and classify all the grid points into interface: those coupled with other subdomain unknowns, and inner: those that are not. Then, the inner-interface ordering arises from applying the mirrored ordering of Fig. 1 to the inner unknowns first and the interface ones afterwards. Remarkably enough, AMGR, the preconditioner derived in Section 3.2, relaxes this constraint, allowing any ordering of the interface unknowns and, most notably, freeing all constraints on the boundary conditions. Hence, AMGR, our final proposal, will not only succeed in outperforming the standard AMG but also apply to a broader range of industrial applications, including, for instance, repeated geometries like the illustrative Fig. 2(b).
@@
where $\bar{K}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{inn}}}$, $\bar{B}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{ifc}}}$, and $\bar{C}\in\mathbb{R}^{n_{\text{ifc}}\times n_{\text{ifc}}}$. Additionally, thanks to the mirrored ordering imposed on the inner and interface grid points, we have that $\bar{K}=\mathbb{I}_{2^{s}}\otimes K$ and $\bar{B}=\mathbb{I}_{2^{s}}\otimes B$. On the other hand, $\bar{C}$ satisfies Eq. 6, and consequently, SpMM can be exploited to accelerate all the submatrix products while reducing their setup costs and memory footprint.
@@ 3.1 Schur complement-based AMG
Given the poor results obtained with LRCAMG($k$), our aim now is to exploit the structure of $\bar{K}$, $\bar{B}$ and $\bar{C}$ to accelerate the AMG approximation of $A$, $M_{A}^{-1}\simeq A^{-1}$. By virtue of Eq. 35, the inverse of $A$’s LDU factorisation reads:
@@
where $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}$ is the Schur complement of $\bar{K}$. Then, we can derive valid preconditioners by seeking approximations of $\bar{K}^{-1}$ and $\bar{S}^{-1}$. In our case, given that $\bar{K}^{-1}=\mathbb{I}_{2^{s}}\otimes K^{-1}$, we will approximate $\bar{K}^{-1}$ through the AMG of $K$, $M_{K}^{-1}\simeq K^{-1}$. This has two major advantages: on the one hand, the setup costs and memory footprint of $M_{K}^{-1}$ are $2^{s}$ times smaller than those of $M_{\bar{K}}^{-1}$. On the other, its application is compatible with SpMM.
Regarding the approximation of $\bar{S}^{-1}$, the better it is, the closer the action of our Schur-based AMG will be to $M_{A}^{-1}$. We have that:
(37) $\bar{S}\coloneqq\bar{C}-\bar{B}^{T}\bar{K}^{-1}\bar{B}=\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}K^{-1}B,$
and $\bar{S}$ containing $K^{-1}$ prevents us from using AMG to approximate $\bar{S}^{-1}$. However, we can obtain a straightforward approximation of $\bar{S}^{-1}$ using FSAI. With this aim, let $L_{K}$ be the lower Cholesky factor of $K$, and $G_{K}\simeq L_{K}^{-1}$ its lower FSAI factor. Then, we have that $G_{K}^{T}G_{K}\simeq K^{-1}$, and we can define:
(38) $\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}G_{K}^{T}G_{K}B\simeq\bar{S},$
whose inverse can be approximated through another FSAI, leading to an approximation that we will denote as $\bar{S}_{G}^{-1}\simeq\bar{S}^{-1}$.
As the model problem will demonstrate, such an approximation of $\bar{S}^{-1}$ is too rough, resulting in slow convergences. Recalling the favourable results of subsection 2.1, we will build an approximation of $\bar{S}^{-1}$ based on LRCFSAI($k$). For simplicity, let us assume a single reflection symmetry. As before, applying a mirrored ordering to the interface grid points ensures that $\bar{C}$ satisfies the following structure:
@@
where $C_{1},C_{2}\in\mathbb{R}^{n_{\text{ifc}}/2\times n_{\text{ifc}}/2}$ account for the interface-interface couplings of the base mesh with itself and its mirroring, respectively. Combining Eqs. 39 and 37, we have that:
(40) $\bar{S}=\begin{pmatrix}S_{1}&C_{2}\\ C_{2}&S_{1}\end{pmatrix},$
@@
After the excellent results obtained with LRCFSAI($k$) and its decoupled corrections, it makes sense to mimic that strategy. First, we will build a rough approximation of $\hat{S}^{-1}$, similar to $\bar{S}_{G}^{-1}$ but based on an FSAI of $S_{1}$ and, therefore, compatible with SpMM. Then, we will correct it through low-rank corrections to make it approach the (block diagonal version of the) most accurate Schur complement at hand:
(42) $\bar{S}_{\text{AMG}}\coloneqq\bar{C}-\mathbb{I}_{2^{s}}\otimes B^{T}M_{K}^{-1}B,$
@@
Similarly to Eq. 38, it is meaningful to approximate $S_{1}^{-1}$ through an FSAI of $C_{1}-B^{T}G_{K}^{T}G_{K}B\simeq S_{1}$:
(43) $G_{S_{1}}^{T}G_{S_{1}}\simeq S_{1}^{-1}.$
Then, by recalling Eq. 41 and the fact that $S_{1}$ is substantially denser than $C_{2}$, we can build the sought approximation of $\hat{S}^{-1}$ as follows:
(44) $\mathbb{I}_{2}\otimes G_{S_{1}}^{T}G_{S_{1}}\simeq\hat{S}^{-1},$
@@
Remarkably enough, to introduce low-rank corrections that improve FSAI, it is not necessary to leverage symmetries and block diagonalise $\bar{S}$. Instead, we could have invoked Theorem 2.1 on $\bar{S}_{G}^{-1}$, correcting it to approach $\bar{S}_{\text{AMG}}$ and giving rise to $\bar{S}_{\text{LRC}}^{-1}\simeq\bar{S}^{-1}$. However, as discussed in Section 2.1 and demonstrated with the model problem, the larger the number of subdomains, the more effective the decoupled corrections.
Regardless of approximating $\bar{S}^{-1}$ with $\bar{S}_{G}^{-1}$, $\bar{S}_{\text{LRC}}^{-1}$, or $Q_{s}^{-1}\hat{S}_{\text{LRC}}^{-1}Q_{s}$, Eq. 36 requires inverting $\bar{K}$ twice, whereas $M_{A}^{-1}$ is only applied once. In order to avoid such an expensive overhead, we can define our Schur-based AMG, henceforth denoted AMGS$(\bar{S}^{-1},\bar{K}^{-1})$, as follows:
(46) $M_{\text{AMGS}}^{-1}(\bar{S}^{-1},\bar{K}^{-1})\coloneqq\begin{pmatrix}\mathbb{I}&-\bar{K}^{-1}\bar{B}\\ &\mathbb{I}\end{pmatrix}\begin{pmatrix}M_{\bar{K}}^{-1}&\\ &\bar{S}^{-1}\end{pmatrix}\begin{pmatrix}\mathbb{I}&\\ -\bar{B}^{T}M_{\bar{K}}^{-1}&\mathbb{I}\end{pmatrix},$
where we compacted the notation of $M_{\bar{K}}^{-1}\coloneqq\mathbb{I}_{2^{s}}\otimes M_{K}^{-1}$ and replaced the extra application of $M_{K}^{-1}$ with $\bar{K}^{-1}$, which, for instance, could be the FSAI of $\bar{K}$, $G_{\bar{K}}^{T}G_{\bar{K}}\coloneqq\mathbb{I}_{2^{s}}\otimes G_{K}^{T}G_{K}$. The main advantages of AMGS$(\bar{S}^{-1},\bar{K}^{-1})$ are being its matrix multiplications compatible with SpMM, together with the smaller memory footprint and setup costs that this allows.
Table 3 summarises the results obtained with AMGS on the model problem of Eq. 25. The results correspond to a MATLAB implementation using an AMG with a PMIS coarsening [18], Extended+I interpolation [17] and an FSAI smoother with 15 nonzeros per row. Conversely, the FSAI of $K$ and of the Schur complement have a pre-filtering tolerance $\tau=0.01$ [30], and 100 and 50 nonzeros per row, respectively. For completeness, three approximations of $\bar{S}^{-1}$ are included in Table 3. Namely, $\bar{S}_{G}^{-1}$, $\bar{S}_{\text{LRC}}^{-1}$, and $Q_{s}^{-1}\hat{S}_{\text{LRC}}^{-1}Q_{s}$. The fact that AMGS$(\bar{S}_{G}^{-1},M_{\bar{K}}^{-1})$ converges almost twice slower than the other AMGS variants confirms the need for low-rank corrections. Especially as $s$ grows, since $n_{\text{ifc}}$ grows accordingly and convergence becomes more sensitive to the approximation of $\bar{S}^{-1}$. Due to this, the convergence of AMGS$(\bar{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$ also deteriorates with $s$. However, thanks to becoming more effective as the number of subdomains grows, decoupled corrections do not suffer from this effect, requiring AMGS$(\hat{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$ roughly the same iterations regardless of $s$. Note that AMGS$(\hat{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$ does actually refer to AMGS$(Q_{s}^{-1}\hat{S}_{\text{LRC}}^{-1}Q_{s},M_{\bar{K}}^{-1})$.
All three variants above require applying $M_{\bar{K}}^{-1}$ twice, which makes the computational costs of the resulting preconditioner exceedingly large. However, as hinted earlier, we can replace the extra application of $M_{\bar{K}}^{-1}$ with FSAI in Eq. 46. Table 3 confirms that we can do this safely, and although it is much lighter, AMGS$(\hat{S}_{\text{LRC}}^{-1},G_{\bar{K}})$ converges comparably to AMGS$(\hat{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$.
Unfortunately, the combination of FSAI and low-rank corrections was not enough to substantially improve the performance of AMGS with respect to LRCAMG($k$) in Table 3. Indeed, it converged more than three times slower than AMG, rendering it impossible to reduce the solution time regardless of the more compute-intensive matrix multiplications. Consequently, we discarded its parallel implementation and developed the final alternative of Section 3.2, which converges comparably to AMG and offers greater flexibility than all previous alternatives.
@@
AMGS$(\bar{S}_{G}^{-1},M_{\bar{K}}^{-1})$ 6 35 38 44
AMGS$(\bar{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$ 6 22 26 29
AMGS$(\hat{S}_{\text{LRC}}^{-1},M_{\bar{K}}^{-1})$ 6 20 23 21
AMGS$(\hat{S}_{\text{LRC}}^{-1},G_{\bar{K}})$ 6 21 23 22
@@ 3.2 Multigrid reduction
As the model problem of Eq. 25 confirmed, the standard AMG algorithm cannot take advantage of spatial symmetries and needs to be appropriately adapted. In particular, we will develop an AMG reduction framework, hereafter denoted as AMGR, that converges comparably to AMG, is compatible with SpMM, and leverages reflection, translational and rotational symmetries. Remarkably enough, AMGR only uses SpMM on the inner unknowns, therefore allowing any interface ordering and, more importantly, freeing all constraints on the boundary conditions, which makes it apply to a significantly broader range of industrial applications without compromising its computational advantages, as $n_{\text{inn}}\gg n_{\text{ifc}}$.
@@
Ideally, to have the fastest possible coarsening, we should only identify as coarse the interface unknowns, i.e., the $n_{\text{ifc}}$ unknowns associated with the interface block, $\bar{C}$. However, realistically sized problems make it impossible to accurately interpolate inner unknowns solely using interface values. This is due to the large connection distance that may occur. Indeed, assuming a $d$-dimensional domain with a similar scale in all directions and $n_{\text{inn}}$ inner unknowns, the maximum inner-interface distance will be of the order of $d]{n_{\text{inn}}}$. Then, a 3D problem with about a million inner unknowns would result in distances of about 100 units, which exceeds the applicability of long-distance interpolation formulas such as Extended+I (ExtI) [17] or Dynamic Pattern Least Squares (DPLS) [43].
@@
Consequently, to allow for an accurate interpolation, we need to convert some inner nodes into coarse. In principle, we can apply any standard coarsening strategy to the inner part of the base mesh. That is, to choose a strength of connection measure, e.g., a classical strength measure, filtering the resulting adjacency graph to obtain $T$, and running a maximum independent set (MIS) algorithm on $T$ to select which inner nodes become coarse. Such a procedure populates the interface with some of the base mesh inner unknowns. Then, to preserve the structure of $\bar{K}=\mathbb{I}_{n_{b}}\otimes K$, we need their analogues in the rest of the subdomains to be set as coarse, too. The problem is that, even if many connections are retained in $T$, the number of coarse nodes generally becomes too large, making the resulting operator’s complexity impractical.
Algorithm 2 summarises the aggressive coarsening strategy designed ad hoc for tackling this problem. Instead of selecting independent nodes from $T$, we select independent nodes from $T^{k}$, for a small power $k$. Using $T^{k}$ is equivalent to constructing an independent set by considering dependent on each other all nodes at a distance smaller or equal to $k$. For simplicity, from now on, $n_{\text{ifc}}$ and $n_{\text{inn}}$ will refer to the populated interface and the remaining inner unknowns. That is, to the number of coarse and fine nodes arising from Algorithm 2, respectively. The same applies to the partitioning of Eq. 35 and $\bar{K}$, $\bar{B}$ and $\bar{C}$.
Algorithm 2 Populated coarsening at a maximum interpolation distance $k$
1:procedure AMGR_Coarsening($A$, $k$)
2:     Measure the strength-of-connection of $A$
3:     Filter the resulting adjacency graph, $T$
4:     Compute the symbolic power of $T$, $T^{k}$
5:     Compute a maximum independent set on $T^{k}$, MIS($T^{k}$)
[…]
7:end procedure
The coarsening of Algorithm 2 yields a prolongation with the following structure:
(47) $P\coloneqq\begin{pmatrix}\bar{W}\\ \mathbb{I}_{n_{\text{ifc}}}\end{pmatrix}\in\mathbb{R}^{n\times n_{\text{ifc}}},$
where $\bar{W}\in\mathbb{R}^{n_{\text{inn}}\times n_{\text{ifc}}}$ and $\mathbb{I}_{n_{\text{ifc}}}\in\mathbb{R}^{n_{\text{ifc}}\times n_{\text{ifc}}}$. Relatively large distances between coarse nodes are not an issue if long-distance interpolations are used. ExtI interpolation is most effective when $k$ is limited to 2. Conversely, DPLS allows for larger values of $k$ and coarser operators. However, there is a trade-off between increasing the value of $k$ to have a more aggressive coarsening and keeping the interpolations accurate. Section 5 investigates the optimal configuration for AMGR, concluding that best results are obtained with $k=2$, DPLS and energy minimisation [31].
In order to make AMGR exploit the computational advantages of SpMM, the top-level smoother is defined as the following block FSAI:
(48) $\begin{pmatrix}G_{\bar{K}}^{T}G_{\bar{K}}&\\ &G_{\bar{C}}^{T}G_{\bar{C}}\end{pmatrix}\simeq A^{-1},$
where $G_{\bar{K}}=\mathbb{I}_{n_{b}}\otimes G_{K}$, $G_{K}^{T}G_{K}\simeq K^{-1}$ and $G_{\bar{C}}^{T}G_{\bar{C}}\simeq\bar{C}^{-1}$. Numerical experiments confirmed that, for relaxation purposes, one can safely ignore $A$’s off-diagonal blocks in Eq. 48, as it causes no significant convergence degradation. Nevertheless, it allows the replacement of SpMV with SpMM and $G_{\bar{K}}$ with $G_{K}$, therefore reducing the memory footprint, setup and application costs of the top-level smoother.
Finally, the reduced operator resulting from the multigrid reduction reads:
(49) $A_{c}\coloneqq P^{T}AP=\bar{W}^{T}\bar{K}\bar{W}+\bar{W}^{T}\bar{B}+\bar{B}^{T}\bar{W}+\bar{C}\in\mathbb{R}^{n_{\text{ifc}}\times n_{\text{ifc}}},$
which is then inverted through a standard AMG. Algorithm 3 summarises the application of AMGR, which aims to enhance the AMG applied in line 5 by reducing the costs of its finest (and most computationally expensive) level. Hence, the multigrid reduction only defines the first level of AMGR, and subsequent levels are purely algebraic.
Algorithm 3 AMGR application in a V$(\nu_{1},\nu_{2})$-cycle
1:procedure AMGR_Apply($A$, $y$, $z$)
2:     Smooth $\nu_{1}$ times $As=y$ starting from $0_{n}$
3:     Compute the residual $r=y-As$
4:     Restrict the residual to the populated interface grid $r_{c}=P^{T}r$
5:     AMG_Apply$\left(A_{c},r_{c},h_{c}\right)$
6:     Prolongate the correction to the entire grid $h=Ph_{c}$
7:     Update $s=s+h$
8:     Smooth $\nu_{2}$ times $Az=y$ starting from $s$
9:end procedure
Table 4 summarises the results obtained with AMGR on the model problem of Eq. 25. The results correspond to a MATLAB implementation of a multigrid reduction using $k=2$, DPLS and energy minimisation. Conversely, the AMG applied to the reduced operator used a PMIS coarsening [18], Extended+I interpolation [17] and an FSAI smoother with 15 nonzeros per row. As it can be seen, adopting the aggressive coarsening that spatial symmetries induce does not harm convergence and allows for the computational advantages of replacing SpMV with SpMM. Hence, given such promising results, we tackled its parallel implementation, which is discussed in Section 4, and evaluated its performance on industrial CFD applications in Section 5.